# Composing Global Optimizers to Reasoning Tasks via Algebraic Objects in Neural Nets

### v3

@@ 5.1 Constructing Partial Solutions with Polynomials
where the generator ${\bm{u}}$ and coefficients ${\bm{c}}_{l}$ are order-1 and the power operation ${\bm{u}}^{l}$ is defined by ring multiplication. The following construction of a polynomial leads to a partial solution.
@@
Suppose ${\bm{u}}$ has 1-set $R_{1}$, $\Omega_{R}({\bm{u}}):=\{r({\bm{u}})|r\in R\}\subseteq\mathbb{C}$ is a set of evaluations on $R$ (multiple values counted once), then if $1\notin\Omega_{R}$, then the polynomial solution $\bm{\rho}_{R}({\bm{u}}):=\prod_{s\in\Omega_{R}({\bm{u}})}({\bm{u}}+\hat{\bm{s}})$ has 0/1-set $(R,R_{1})$ up to a scale. Here $\hat{\bm{s}}$ is any order-1 weight that satisfies $r(\hat{\bm{s}})=-s$ for any $r\in R\cup R_{+}$. For example, $\hat{\bm{s}}=-s^{1/3}{\bm{1}}$.
@@
For convenience, we use $\bm{\rho}({\bm{u}})$ to represent the maximal polynomial, i.e., when $R=\arg\max_{1\notin\Omega_{R}({\bm{u}})}|\Omega_{R}({\bm{u}})|$ is the largest subset of MPs with $1\notin\Omega_{R}({\bm{u}})$. Our goal is to find low-order (partial) solutions, since gradient descent prefers low order solutions (see Theorem 6). Although there exist high-degree but low-order polynomials, e.g., ${\bm{u}}^{9}+{\bm{1}}$, in general, degree $L$ and order $q$ are correlated, and we can find low-degree ones instead. To achieve that, ${\bm{u}}$ should be properly selected (e.g., symmetric weights) to create as many duplicate values (but not $1$) in $R$ as possible.
Evaluation on MPs
$R_{\mathrm{c}}$ $R_{\mathrm{n}}$ $R_{*}$ Maximal
Symbol $[a,b,c]$ $\bar{a}bc$ $a\bar{b}c$ $ab\bar{c}$ $\bar{a}ac$ $\bar{b}bc$ $aac$ $bbc$ $\bar{a}\bar{a}c$ $\bar{b}\bar{b}c$ polynomial $\bm{\rho}({\bm{u}})$ order $q$
${\bm{1}}_{k}$ $[1,1,1]$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
$\tilde{\bm{1}}_{k}$ $[-1,-1,1]$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
${\bm{u}}_{\mathrm{one}}$ $[1,-1,-1]$ $1$ $1$ $1$ $-1$ $-1$ $-1$ $-1$ $-1$ $-1$ ${\bm{u}}+{\bm{1}}$ 2
${\bm{u}}_{\mathrm{syn}}$ $[\omega_{3},\omega_{3},\omega_{3}]$ $\omega_{3}$ $\omega_{3}$ $\omega_{3}$ $\omega_{3}$ $\omega_{3}$ $1$ $1$ $\bar{\omega}_{3}$ $\bar{\omega}_{3}$ ${\bm{u}}^{2}+{\bm{u}}+{\bm{1}}$ 3
${\bm{u}}_{3\mathrm{c}}$ $[\omega_{3},\bar{\omega}_{3},1]$ $\omega_{3}$ $\bar{\omega}_{3}$ $1$ $1$ $1$ $\bar{\omega}_{3}$ $\omega_{3}$ $\omega_{3}$ $\bar{\omega}_{3}$ ${\bm{u}}^{2}+{\bm{u}}+{\bm{1}}$ 3
${\bm{u}}_{3\mathrm{a}}$ $[1,\omega_{3},\bar{\omega}_{3}]$ $1$ $\omega_{3}$ $\bar{\omega}_{3}$ $\bar{\omega}_{3}$ $\bar{\omega}_{3}$ $\bar{\omega}_{3}$ $\omega_{3}$ $\bar{\omega}_{3}$ $1$ ${\bm{u}}^{2}+{\bm{u}}+{\bm{1}}$ 3
${\bm{u}}_{4\mathrm{c}}$ $[\mathrm{i},-\mathrm{i},1]$ $-1$ $-1$ $1$ $1$ $1$ $-1$ $-1$ $-1$ $-1$ ${\bm{u}}+{\bm{1}}$ 2
${\bm{u}}_{4\mathrm{a}}$ $[1,\mathrm{i},-\mathrm{i}]$ $1$ $-1$ $-1$ $-\mathrm{i}$ $-\mathrm{i}$ $-\mathrm{i}$ $\mathrm{i}$ $-\mathrm{i}$ $\mathrm{i}$ ${\bm{u}}^{3}+{\bm{u}}^{2}+{\bm{u}}+{\bm{1}}$ 4
${\bm{u}}_{\nu}$ $[\nu,-\nu,-\bar{\nu}^{2}]$ $\nu^{2}$ $\nu^{2}$ $\nu^{4}$ $-\bar{\nu}^{2}$ $-\bar{\nu}^{2}$ $-1$ $-1$ $-\nu^{4}$ $-\nu^{4}$ 9-th degree 10
@@
Table 1: Exemplar order-1 single frequency generator ${\bm{u}}^{(k)}$ with $r_{kkk}({\bm{u}}^{(k)})=1$. In the single-frequency case, for each MP $r$ we use “$\bar{a}bc$” to represent $r_{-k,k,k}$ and “$\bar{a}\bar{a}c$” to represent $r_{a,-k,-k,k}$, etc. We omit superscript “${(k)}$” for clarity and omit conjugate columns (i.e., $\bar{a}\bar{b}c$ which is conjugate to $ab\bar{c}$). Here, $\omega_{3}:=e^{2\pi\mathrm{i}/3}$ and $\omega_{4}:=\mathrm{i}$ are the 3rd and 4th roots of unity. The constructed solutions are partial, i.e., the evaluation of some MPs yields $1$ (red cell) and cannot be the root of the polynomial according to Theorem 4. Note that ${\bm{u}}_{\nu}$ is a general case with ${\bm{u}}_{\nu=1}={\bm{u}}_{\mathrm{one}}$ and ${\bm{u}}_{\nu=\mathrm{i}}={\bm{u}}_{4\mathrm{c}}$.
5.2 Composing Global Solutions
@@
We first consider the case that the generator ${\bm{u}}$ is only nonzero at frequency $k$ (and thus $-k$ by Hermitian constraints), but zero in other frequencies, i.e., $u_{pk^{\prime}0}=0$ for $k^{\prime}\neq\pm k$. Such solutions correspond to Fourier bases in the original domain. Also, ${\bm{u}}$ has 1-set $R_{1}=\{r_{kkk}\}$. This means that ${\bm{u}}$ can be characterized by three numbers $u_{ak0}=a$, $u_{bk0}=b$, and $u_{ck0}=c$ with $abc=1$. In this case, only a subset of monomial potentials (MPs) whose indices only involve a single frequency $k$ are non-zero (e.g., $r_{k,-k,k}\in R_{\mathrm{c}}$ and $r_{b,-k,k,k}\in R_{\mathrm{n}}$), which makes our construction much easier.
@@
Following Theorem 4, we can construct different partial solutions. Some examples are shown in Table 1, which do not reach the complete set $R_{\mathrm{c}}\cup R_{\mathrm{n}}\cup R_{*}$ and therefore are not global. Note that it is possible to create a generator so that all MPs are not $1$ (e.g., ${\bm{u}}_{3\mathrm{c}}*{\bm{u}}_{4\mathrm{a}}$), but then $|\Omega_{R}({\bm{u}})|$ will be too large, producing high-degree polynomials (e.g., ${\bm{u}}_{3\mathrm{c}}*{\bm{u}}_{4\mathrm{a}}$ gives a 10-th-degree polynomial).
However, utilizing these partial solutions, with Lemma 2 we can construct global optimizers:
Corollary 2 (Order-6 global optimizers).
@@
The following ``$3\times 2"$ Fourier solutions satisfy the sufficient condition (Lemma 1) and thus are global optimizers (assuming $d$ is odd):
@@
Other solutions. We may replace ${\bm{u}}_{\mathrm{syn}}$ and ${\bm{u}}_{\nu}$ with any other pairs that collectively cover all MPs. For example, ${\bm{u}}_{\mathrm{syn}}$ can be combined with any of $\{{\bm{u}}_{3\mathrm{c}},{\bm{u}}_{3\mathrm{a}},{\bm{u}}_{4\mathrm{a}}\}$, and ${\bm{u}}_{\nu=\pm\mathrm{i}}$ can be coupled with ${\bm{u}}_{3\mathrm{a}}$ or ${\bm{u}}_{4\mathrm{a}}$, etc. Here we pick one with a small order. Compared to construction from Gromov (2023), ours is much more concise and does not use infinite-width approximation.
Figure 3: Solutions obtained by the Adam optimizers on $\ell_{2}$ loss for modular addition task with $|G|=d=7$ and $q=20$ hidden nodes. Top: For each frequency ±$\pm k$, there are exactly $6$ hidden nodes representing such a frequency, consistent with Corollary 1. Bottom: Optimizing Eqn. 3 without the last term $\sum_{m\neq 0}\sum_{p\in\{a,b\}}\Big{|}\sum_{k^{\prime}}r_{p,k^{\prime},m-k^{\prime},k}\Big{|}^{2}$ (equivalently removing the constraint $R_{*}$). Now each frequency has exactly $3$ hidden nodes, which corresponds to the solution ${\bm{z}}_{\mathrm{syn}}=\bm{\rho}({\bm{u}}_{\mathrm{syn}})$ in Tbl. 1.
Even $d$. For even $d$, simply replace $(d-1)/2$ with $\lfloor(d-1)/2\rfloor$ and add an additional order-2 term $\bm{\rho}({\bm{u}}_{\mathrm{one}})={\bm{u}}_{\mathrm{one}}+{\bm{1}}$ (Tbl. 1) for the frequency $d/2$. Note that the frequency $k=d/2$ only has $r_{kkk}$, $r_{akkk}$ and $r_{bkkk}$, and all other conjugate combinations are absent. Thus ${\bm{u}}^{(k)}_{\mathrm{one}}+{\bm{1}}_{k}$ covers them all.
Fig. 3 shows a case with $d=7$. In this case, each frequency, out of $(d-1)/2=3$ total number of frequencies, is associated with $6$ hidden nodes. If we remove the last term in the loss that corresponds to $R_{*}$, then an order-3 solution suffices (i.e. ${\bm{z}}_{\mathrm{syn}}=\bm{\rho}({\bm{u}}_{\mathrm{syn}})$).
Using polynomials, we can also construct perfect-memorization solutions. For this, we first define two generators ${\bm{u}}_{a}$ with $u^{(\alpha)}_{\cdot k0}=[\omega_{d}^{k},1,\bar{\omega}_{d}^{k}]\mathbb{I}(k\neq 0)$, and ${\bm{u}}_{b}$ with $u^{(\beta)}_{\cdot k0}=[1,\omega_{d}^{k},\bar{\omega}_{d}^{k}]\mathbb{I}(k\neq 0)$. Here $\omega_{d}:=e^{2\pi\mathrm{i}/d}$ is the $d$-th root of unity.
@@
We construct two $d$-order weights ${\bm{z}}_{a}$ and ${\bm{z}}_{b}$:
${\bm{z}}_{a}=\sum_{j=0}^{d-1}{\bm{u}}_{a}^{j},\quad\quad{\bm{z}}_{b}=\sum_{j=0}^{d-1}{\bm{u}}_{b}^{j}$ (7)
Here ${\bm{z}}_{a}\in R_{\mathrm{c}}(k_{1}\neq k)\cap R_{\mathrm{n}}\cap R_{*}(p=b\mathrm{\ or\ }m\neq k)$, ${\bm{z}}_{b}\in R_{\mathrm{c}}(k_{2}\neq k)\cap R_{\mathrm{n}}\cap R_{*}(p=a\mathrm{\ or\ }m\neq k)$. Then ${\bm{z}}_{M}=d^{-2/3}{\bm{z}}_{a}*{\bm{z}}_{b}$ satisfies the sufficient condition (Lemma 1) and is the perfect memorization solution with $\mathrm{ord}({\bm{z}}_{M})=d^{2}$:
[…]
@@
To see why this corresponds to perfect memorization, simply apply an inverse Fourier transform for each hidden node $(j_{1},j_{2})$, and the original weights are (zero-mean) delta function located at $j_{1}$, $j_{2}$ and $j_{1}+j_{2}$ accordingly.
Interestingly, there also exists a lower-order solution, $2\times 2$, that meets $R_{\mathrm{c}}$ and $R_{*}$ but not $R_{\mathrm{n}}$:
@@
While ${\bm{z}}^{(k)}_{F4}$ itself does not satisfy the sufficient condition (Eqn. 4), it is part of a global optimizer when mixing with ${\bm{z}}_{F6}$:
Corollary 5 (Mixed order-4/6 global optimizers).
With ${\bm{z}}^{(k)}_{F4}$, there is a global optimizer to Eqn. 3 that does not meet the sufficient condition, i.e., $\sum_{k^{\prime}}r_{p,k^{\prime},-k^{\prime},m}=0$ but $r_{p,k^{\prime},-k^{\prime},m}\neq 0$:
@@
Remarks. To construct $\hat{\bm{z}}_{F6}$, in addition to ${\bm{z}}_{\mathrm{syn}}*{\bm{z}}_{\nu=1}$, we could use other compositions of single frequency solutions to achieve the same effects. For example, ${\bm{z}}_{\mathrm{syn},\alpha\beta}*{\bm{z}}_{\nu=\mathrm{i}}$, where ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ is:
$z_{ak\cdot}=[1,\omega_{3}\alpha,\bar{\omega}_{3}\beta],\quad\quad z_{bk\cdot}=[1,\omega_{3}\bar{\alpha},\bar{\omega}_{3}\bar{\beta}],\quad\quad z_{ck\cdot}=[1,\omega_{3},\bar{\omega}_{3}]$ (11)
Note that multiple per frequency order-6 solutions can be inserted in this construction. Compared to all order-6 solutions ${\bm{z}}_{F6}$, this ${\bm{z}}_{F4/6}$ mixture solution has a lower order and is perceived in the experiments (See Fig. 7), in particular when $d$ is large (Tbl. 2), showing a strong preference of gradient descent towards lower order solutions.
Necessary Conditions. For now we have discussed solutions constructed to be global optimizers (i.e. sufficient conditions), For the structure that weights must follow to satisfy Eqn. 1 (i.e. necessary conditions), please check Lemma 5 and Lemma 6 in Appendix Sec. C.3.
Figure 4: Dynamics of monomial potentials (MPs) over the training process for modular addition with $d=23$ and $q=1024$ hidden nodes. Top Row. Left: Training/test accuracy reaches 100% and loss close to 00. Test accuracy jumps after training reaches 100% (grokking). Mid: After 10k epochs, the distribution of solution orders are concentrated at 4 and 6 (Corollary 2 and 4). Right: Dynamics of $r_{k_{1}k_{2}k}$. Summation of diagonal $r_{kkk}$ converges towards $d-1$ (dotted line) with ripple effects, while off-diagonal $r_{k_{1}k_{2}k}$ converges towards 00. Bottom Row. Dynamics of different MPs. Order-4 and order-6 behave differently on $r_{p,k,-k,k}$, because order-4 does not satisfy the sufficient condition (Lemma 1) but a mixture of order-4 and order-6 (i.e., ${\bm{z}}_{F4/6}$) is still the global optimizer to the $L_{2}$ loss (Corollary 5).
6 Gradient dynamics
Now we have characterized the structures of global optimizers. One natural question arises: why does the optimization procedure not converge to the perfect memorization solution ${\bm{z}}_{M}$, but to the Fourier solutions ${\bm{z}}_{F6}$ and ${\bm{z}}_{F4/6}$? The answer is given by gradient dynamics.
Let ${\bm{r}}=[r_{k_{1}k_{2}k},r_{pk_{1}k_{2}k}]\in\mathbb{C}^{4d^{3}}$ be a vector of all MPs, and $J:=\frac{\partial{\bm{r}}}{\partial{\bm{z}}}\frac{\partial{\bm{z}}}{\partial\mathcal{W}}$ be the Jacobian matrix of the mapping ${\bm{r}}={\bm{r}}({\bm{z}}(\mathcal{W}))$ in which $\mathcal{W}$ is the collection of original weights. Note that when we take derivatives with respect to $r$ and apply chain rules, we treat $r$ and its complex conjugate (e.g., $r_{kkk}$ and $r_{-k,-k,-k}=\bar{r}_{kkk}$) as independent variables. Since we run the gradient descent on $\mathcal{W}$, will such (indirect) optimization leads to a descent of ${\bm{r}}$ towards the desired targets (Lemma 1)? This is confirmed by the following theorem:
Theorem 5 (Dynamics of MPs).
The dynamics of MPs satisfies $\dot{\bm{r}}=-JJ^{*}\overline{\nabla_{\bm{r}}\ell}$, which has positive inner product with the negative gradient direction $-\overline{\nabla_{\bm{r}}\ell}$.
Figure 5: Solution distribution over different weight decay regularization for $q=512$, trained with 10k epochs with Adams with learning rate $0.01$ on modular addition (i.e., predicting $a+b\mod d$) with $d\in\{23,71,127\}$. The two red dashed lines correspond to order-4/6 solutions. The histogram is accumulated over 5 random seeds.
Corollary 1 shows that by ring multiplication, we could create infinitely many global optima from a base one. The following theorem answers which solution gradient dynamics picks.

### v4

@@ 5 Composing Global Solutions
where the generator $\bm{u}$} and coefficients ${\bm{c}}_{l}$ are order-1 and the power operation ${\bm{u}}^{l}$ is defined by ring multiplication. Then a partial solution can be constructed:
@@
[…]
@@
For convenience, we use $\bm{\rho}({\bm{u}})$ to represent the maximal polynomial, i.e., when $R=\arg\max_{1\notin\Omega_{R}({\bm{u}})}|\Omega_{R}({\bm{u}})|$ is the largest subset of SPs with $1\notin\Omega_{R}({\bm{u}})$. Our goal is to find low-order (partial) solutions, since gradient descent prefers low order solutions (see Theorem 5). Although there exist high-degree but low-order polynomials, e.g., ${\bm{u}}^{9}+{\bm{1}}$, in general, degree $L$ and order $q$ are correlated, and we can find low-degree ones instead. To achieve that, $\bm{u}$} should be properly selected (e.g., symmetric weights) to create as many duplicate values (but not $1$) in $R$ as possible.
Evaluation on SPs
$R_{\mathrm{c}}$ $_{\mathrm{n}} RR_{*}$ Maximal
[…]
${\bm{1}}_{k}$ $[1,1,1]$ 11 11 11 11 11 11 11 11 11
$\tilde{\bm{1}}_{k}$ [1,1,1][-1,-1,1] 11 11 11 11 11 11 11 11 11
${\bm{u}}_{\mathrm{one}}$ [1,1,1][1,-1,-1] 11 11 11 1-1 1-1 1-1 1-1 1-1 1-1 ${\bm{u}}+{\bm{1}}$ 2
${\bm{u}}_{\mathrm{syn}}$ $\omega_{3},\omega_{3},\omega_{3}]$ $\omega_{3}$ $\omega_{3}$ $\omega_{3}$ $\omega_{3}$ $\omega_{3}$ 11 11 $\bar{\omega}_{3}$ $\bar{\omega}_{3}$ ${\bm{u}}^{2}+{\bm{u}}+{\bm{1}}$ 3
${\bm{u}}_{3\mathrm{c}}$ $\omega_{3},\bar{\omega}_{3},1]$ $\omega_{3}$ $\bar{\omega}_{3}$ 11 11 11 $\bar{\omega}_{3}$ $\omega_{3}$ $\omega_{3}$ $\bar{\omega}_{3}$ ${\bm{u}}^{2}+{\bm{u}}+{\bm{1}}$ 3
${\bm{u}}_{3\mathrm{a}}$ $\omega_{3},\bar{\omega}_{3}] 11$ $\omega_{3}$ $\bar{\omega}_{3}$ $\bar{\omega}_{3}$ $\bar{\omega}_{3}$ $\bar{\omega}_{3}$ $\omega_{3}$ $\bar{\omega}_{3}$ 11 ${\bm{u}}^{2}+{\bm{u}}+{\bm{1}}$ 3
${\bm{u}}_{4\mathrm{c}}$ $\mathrm{i},-\mathrm{i},1]$ 1-1 1-1 11 11 11 1-1 1-1 1-1 1-1 ${\bm{u}}+{\bm{1}}$ 2
${\bm{u}}_{4\mathrm{a}}$ $\mathrm{i},-\mathrm{i}] 11$ 1-1 1-1 $\mathrm{i}$ $\mathrm{i}$ $\mathrm{i}$ $\mathrm{i}$ $\mathrm{i}$ $\mathrm{i}$ ${\bm{u}}^{3}+{\bm{u}}^{2}+{\bm{u}}+{\bm{1}}$ 4
${\bm{u}}_{\nu}$ $\nu,-\nu,-\bar{\nu}^{2}]$ $\nu^{2}$ $\nu^{2}$ $\nu^{4}$ $\bar{\nu}^{2}$ $\bar{\nu}^{2}$ 1-1 1-1 $\nu^{4}$ $\nu^{4}$ 9-th degree 10
@@
Table 1: Exemplar order-1 single frequency generator ${\bm{u}}^{(k)}$ with $r_{kkk}({\bm{u}}^{(k)})=1$. In the single-frequency case, for each MP $r$ we use $\bar{a}bc$” to represent $r_{-k,k,k}$ and $\bar{a}\bar{a}c$” to represent $r_{a,-k,-k,k}$, etc. For brevity, superscript “(k){(k)}” and conjugate columns (i.e., $\bar{a}\bar{b}c$ conjugate to $ab\bar{c}$) are omitted. Here, $\omega_{3}:=e^{2\pi\mathrm{i}/3}$ and $\omega_{4}:=\mathrm{i}$ are 3rd/4th roots of unity. The constructions are partial, i.e., the evaluation of some SPs yields 11 (red cell) and cannot be the root of the polynomial (Theorem 4). Note that ${\bm{u}}_{\nu}$ is a general case with ${\bm{u}}_{\nu=1}={\bm{u}}_{\mathrm{one}}$ and ${\bm{u}}_{\nu=\mathrm{i}}={\bm{u}}_{4\mathrm{c}}$.
@@
Composing Global Solutions. We first consider the case that the generator $\bm{u}$} is only nonzero at frequency $k$ (and thus k-k by Hermitian constraints), but zero in other frequencies, i.e., $u_{pk^{\prime}0}=0$ for $k^{\prime}\neq\pm k$. Such solutions correspond to Fourier bases in the original domain. Also, $\bm{u}$} has 1-set $R_{1}=\{r_{kkk}\}$. This means that $\bm{u}$} can be characterized by three numbers $u_{ak0},u_{bk0},u_{ck0}$ with $u_{ak0}u_{bk0}u_{ck0}=1$. In this case, only a subset of sum potentials (SPs) whose indices only involve a single frequency $k$ are non-zero (e.g., $r_{k,-k,k}\in R_{\mathrm{c}}$ and $r_{b,-k,k,k}\in R_{\mathrm{n}}$), facilitating our construction.
@@
Following Theorem 4, we can construct different partial solutions. Some examples are shown in Table 1. These solutions do not make all sum potentials in $R_{\mathrm{c}}\cup R_{\mathrm{n}}\cup R_{*}$ vanish and therefore are not global. Note that it is possible to create a global solution this way, but then $|\Omega_{R}({\bm{u}})|$ will be too large, producing high-degree/order polynomials (e.g., ${\bm{u}}_{3\mathrm{c}}*{\bm{u}}_{4\mathrm{a}}$ gives a 10th-degree polynomial). Instead, utilizing these partial solutions, with Lemma 2 we can construct global solutions with smaller orders:
Corollary 2 (Order-6 global solutions).
@@
The following $3\times 2$" Fourier solutions satisfy the sufficient condition (Lemma 1) and thus are global solutions when $d$ is odd:
@@
Remarks. We may replace ${\bm{u}}_{\mathrm{syn}}$ and ${\bm{u}}_{\nu}$ with other pairs that collectively cover all SPs. For example, ${\bm{u}}_{\mathrm{syn}}$ can be combined with any of $\{{\bm{u}}_{3\mathrm{c}},{\bm{u}}_{3\mathrm{a}},{\bm{u}}_{4\mathrm{a}}\}$, and ${\bm{u}}_{\nu=\pm\mathrm{i}}$ can be coupled with ${\bm{u}}_{3\mathrm{a}}$ or ${\bm{u}}_{4\mathrm{a}}$, etc. Here we pick one with a small order. Compared to Gromov (2023), our construction is more concise without infinite-width approximation. For even $d$, simply replace (d1)/2(d-1)/2 with $\lfloor(d-1)/2\rfloor$ and add an additional order-2 term $\bm{\rho}({\bm{u}}_{\mathrm{one}})={\bm{u}}_{\mathrm{one}}+{\bm{1}}$ (Tbl. 1) for frequency $k=d/2$, which only has $r_{kkk}$, $r_{akkk}$ and $r_{bkkk}$, and all other combinations are absent.
Figure 2: Solutions obtained by Adam optimizers on $\ell_{2}$ loss for modular addition task with |G|=d=7|G|\!=\!d\!=\!7 and q=20q\!=\!20 hidden nodes. Top: For each frequency $\pm k$, exactly 66 hidden nodes exist (Corollary 2). Bottom: Optimizing Eqn. 3 without the last term $\sum_{m\neq 0}\sum_{p\in\{a,b\}}\Big|\sum_{k^{\prime}}r_{p,k^{\prime},m-k^{\prime},k}\Big|^{2}$ (i.e., without constraint $R_{*}$). Now each frequency has exactly 33 hidden nodes, corresponding to the solution ${\bm{z}}_{\mathrm{syn}}=\bm{\rho}({\bm{u}}_{\mathrm{syn}})$ in Tbl. 1.
Fig. 2 shows a case with $d=7$. In this case, each frequency, out of (d1)/2=3(d-1)/2=3 total number of frequencies, is associated with $6$ hidden nodes. If we remove the last term in the loss that corresponds to $R_{*}$, then an order-3 solution suffices (i.e. ${\bm{z}}_{\mathrm{syn}}=\bm{\rho}({\bm{u}}_{\mathrm{syn}})$). Perfect-memorization solutions can also be constructed. Let two generators be ${\bm{u}}_{\alpha}$ with $u^{(\alpha)}_{\cdot k0}\!=\![\omega_{d}^{k},1,\bar{\omega}_{d}^{k}]\mathbb{I}(k\!\neq\!0)$, and ${\bm{u}}_{\beta}$ with $u^{(\beta)}_{\cdot k0}\!=\![1,\omega_{d}^{k},\bar{\omega}_{d}^{k}]\mathbb{I}(k\!\neq\!0)$. Here $\omega_{d}:=e^{2\pi\mathrm{i}/d}$ is the $d$-th root of unity. Then,
@@
We construct two $d$-order weights $({\bm{z}}_{\alpha},{\bm{z}}_{\beta})$ from order-1 generators $({\bm{u}}_{\alpha},{\bm{u}}_{\beta})$:
${\bm{z}}_{\alpha}=\sum_{j=0}^{d-1}{\bm{u}}_{\alpha}^{j},\quad\quad{\bm{z}}_{\beta}=\sum_{j=0}^{d-1}{\bm{u}}_{\beta}^{j}$ (7)
Here ${\bm{z}}_{\alpha}\in R_{\mathrm{c}}(k_{1}\neq k)\cap R_{\mathrm{n}}\cap R_{*}(p=b\mathrm{\ or\ }m\neq k)$, ${\bm{z}}_{\beta}\in R_{\mathrm{c}}(k_{2}\neq k)\cap R_{\mathrm{n}}\cap R_{*}(p=a\mathrm{\ or\ }m\neq k)$. Then ${\bm{z}}_{M}=d^{-2/3}{\bm{z}}_{\alpha}*{\bm{z}}_{\beta}$ satisfies the sufficient condition (Lemma 1) and is the perfect memorization solution with $\mathrm{ord}({\bm{z}}_{M})=d^{2}$:
$z^{(M)}_{akj_{1}j_{2}}=d^{-\frac{2}{3}}\omega^{kj_{1}},\quad z^{(M)}_{bkj_{1}j_{2}}=d^{-\frac{2}{3}}\omega^{kj_{2}},\quad z^{(M)}_{ckj_{1}j_{2}}=d^{-\frac{2}{3}}\omega^{-k(j_{1}+j_{2})}$ (8)
@@
To see why this corresponds to perfect memorization, simply apply an inverse Fourier transform for each hidden node $(j_{1},j_{2})$, which leads to (zero-mean) delta weights located at $j_{1}$, $j_{2}$ and $j_{1}+j_{2}$. Interestingly, there also exists a lower-order solution, $2\times 2$, that meets $R_{\mathrm{c}}$ and $R_{*}$ but not $R_{\mathrm{n}}$:
@@
Although ${\bm{z}}^{(k)}_{F4}$ does not satisfy the sufficient condition (Eqn. 4), it is part of a global solution when mixed with ${\bm{z}}_{F6}$:
Corollary 5 (Mixed order-4/6 global solutions).
With ${\bm{z}}^{(k)}_{F4}$, there is a global solution to Eqn. 3 that does not meet the sufficient condition, i.e., $\sum_{k^{\prime}}r_{p,k^{\prime},-k^{\prime},m}=0$ but $r_{p,k^{\prime},-k^{\prime},m}\neq 0$:
@@
The specific formats of $\hat{\bm{z}}^{(k_{0})}_{F6}$ is shown in Appendix (please check the proof and Eqn. 68). Multiple order-6 solutions per frequency can be inserted in this construction. Compared to ${\bm{z}}_{F6}$, this order-4/6 mixture solution has a lower order and is perceived in the experiments (See Fig. 6), in particular when $d$ is large (Tbl. 2), showing a strong preference of gradient descent towards lower order solutions.
Figure 3: Dynamics of sum potentials (SPs) over the training process for modular addition with $d=23$ and $q=1024$ hidden nodes. Top Row. Left: Training/test accuracy reaches 100% and loss close to 0. Test accuracy jumps after training reaches 100% (grokking). Mid: After 10k epochs, the distribution of solution orders are concentrated at 4 and 6 (Corollary 2 and 4). Right: Dynamics of $r_{k_{1}k_{2}k}$. Summation of diagonal $r_{kkk}$ converges towards d1d-1 (dotted line) with ripple effects, while off-diagonal $r_{k_{1}k_{2}k}$ converges towards 0. Bottom Row. Dynamics of different SPs. Order-4 and order-6 behave differently on $r_{p,k,-k,k}$, because order-4 does not satisfy the sufficient condition (Lemma 1) but a mixture of order-4 and order-6 (i.e., ${\bm{z}}_{F4/6}$) is still the global solution to the $L_{2}$ loss (Corollary 5).
6 Exploring the solution solution with Gradient dynamics
Now we have characterized the structures of global solutions. One natural question arises: why does the optimization procedure not converge to the perfect memorization solution ${\bm{z}}_{M}$, but to the Fourier solutions ${\bm{z}}_{F6}$ and ${\bm{z}}_{F4/6}$? Although characterizing the full gradient dynamics is beyond the scope of this paper, we theoretically characterize some rough behaviors below.
Corollary 1 shows that by ring multiplication, we could create infinitely many global solutions from one. Then Thm. 5 answers which solution the gradient dynamics may pick: